OpenStudy (anonymous):

How many moles are equivalent to 2.50x1020 atoms of Fe?

OpenStudy (chmvijay):

6.023*23 atoms of Fe = 1mole 2.50*10^20 atoms of Fe =2.50*10^20 / 6.023*10^23 =can you find it

OpenStudy (anonymous):

Preciate it man !